OpenStudy (anonymous):

Differentiate the function. g(t) = 4t^−1/8

OpenStudy (kmeis002):

Power rule states \[ \frac{d}{dx} (kx^n) = knx^{n-1}\] Where k and n are Real constants. So what is k and n in your case?

OpenStudy (anonymous):

uh... I'm going to say -1/8?

OpenStudy (anonymous):

and 4?

OpenStudy (anonymous):

see what i think you would do to find the answer is multiply 4 times (1/8) and then subtract 1 from the exponent

OpenStudy (kmeis002):

That is correct. Put all those things together and you have your answer

OpenStudy (anonymous):

but when i put them together the answer doesn't make sense

OpenStudy (anonymous):

i get -.6250538953 which can't be right

OpenStudy (kmeis002):

One moment.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[g(t) = 4t^{−1/8}\]\[g'(x)=-\frac{1}{8}\times 4x^{-\frac{1}{8}-1}\] is a start

OpenStudy (anonymous):

then \[g'(x)=-\frac{1}{2}x^{-\frac{9}{8}}\]

OpenStudy (anonymous):

power rule right?

OpenStudy (anonymous):

when i plug that into my calc is says domain error

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait how did you get 9/8?

OpenStudy (anonymous):

the fractions are confusing me

OpenStudy (anonymous):

like i said in the last post the hardest part of calc one is adding and subtracting fractions

OpenStudy (anonymous):

when i plug the answer you gave me on my online homework it says its incorrect

OpenStudy (anonymous):

btw i got \(-\frac{9}{8}\) \[-\frac{1}{8}-1=-\frac{1}{8}-\frac{8}{8}=-\frac{9}{8}\]

OpenStudy (anonymous):

did you plug in 9/8 or -9/8 for the exponent ?

OpenStudy (anonymous):

-9/8

OpenStudy (anonymous):

and this is the function right? \[g(t) = 4t^{−1/8}\]

OpenStudy (anonymous):

\[4t ^{-1/8}\]

OpenStudy (anonymous):

is the equation

OpenStudy (anonymous):

then i will stick with \[g'(t)=-\frac{1}{2}t^{-\frac{9}{8}}\]

OpenStudy (anonymous):

I'm got the same exact answer i don't understand why its not working. this is why i hate online math!

OpenStudy (kmeis002):

Can you take a screenshot of how you entered it?

OpenStudy (anonymous):

sometimes i am wrong, but this one i would bet \(\$1,000\) on make sure to use a \(t\) as the variable, maybe that was the mistake

OpenStudy (anonymous):

how do you take a screenshot on a macbook pro?

OpenStudy (anonymous):

there are many reasons to hate on line math that is for sure webassign? mymathlabs?

OpenStudy (anonymous):

printscreen button i think says maybe prtscr

OpenStudy (anonymous):

webassign!

OpenStudy (kmeis002):

I believe its (⌘)-Shift-3. They should be saved on yoru desktop

OpenStudy (anonymous):

Press Command (⌘)-Shift-3.

OpenStudy (kmeis002):

Thats the mac button, the clover + shift-3

OpenStudy (anonymous):

use a \(t\) not an \(x\)

OpenStudy (anonymous):

omg wow! thank you.

OpenStudy (anonymous):

im so sorry!

OpenStudy (anonymous):

no it was my fault

OpenStudy (anonymous):

i inadvertently switches to x's cause that is what i am used to

OpenStudy (anonymous):

calc one on line?

OpenStudy (anonymous):

no i would of done the same i tried it before like that and it didn't work so don't worry about it and yeah the homework is.

OpenStudy (anonymous):

mind helping with one more? I'm sure you do the same with the exponent and everything \[f(x)=\frac{ 2 }{ 5 }x ^{10}\]

OpenStudy (anonymous):

same damned thing

OpenStudy (anonymous):

\[f'(x)=4x^9\] now it is easier, not fractions

OpenStudy (anonymous):

i'm thinking you multiply the fraction and subtract 1 from the exponent

OpenStudy (anonymous):

yeah \[10\times \frac{2}{5}=4\]

OpenStudy (anonymous):

okay i just wanted to make sure i was right so it wouldn't lock me out!

OpenStudy (anonymous):

lol good more?

OpenStudy (anonymous):

i have a LOT lol do you know anything about differentiate

OpenStudy (anonymous):

Differentiate \[g(x)= 3e ^{x}\sqrt{x}\]

OpenStudy (anonymous):

that is what you are doing differentiating

OpenStudy (anonymous):

I know x square is equivalent to x1/2

OpenStudy (anonymous):

product rule for this one \[(fg)'-f'g+g'f\] with \[f(x)=e^x,f'(x)=e^x,g(x)=\sqrt{x}, g'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

typo there are on the first line, i meant \[(fg)'=f'g+g'f\]

OpenStudy (anonymous):

yeah you are right about x^2 but let me give you a small piece of advice

OpenStudy (anonymous):

okay whats the advice lol, i need all i can get

OpenStudy (anonymous):

the square root is a very very common function math teachers love to put it on quizzes and exams while your colleages are differentiating it by writing \[\sqrt{x}=x^{\frac{1}{2}}\\ \frac{d}{dx}[x^{\frac{1}{2}}]=\frac{1}{2}x^{\frac{1}{2}-1}\] etc, just memorize it , it never changes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[\frac{d}{dx}[\sqrt{x}]=\frac{1}{2\sqrt{x}}\] and that is all like knowing \(7\times 8=56\) just know it

OpenStudy (anonymous):

okay so then what is our first step in solving this problem

OpenStudy (anonymous):

knowing that it is a product and that you have to use the product rule

OpenStudy (anonymous):

\[\large (fg)'=f'g+g'f\]

OpenStudy (anonymous):

so we multiply it by the same thing

OpenStudy (anonymous):

in this case \[f(x)=e^x,f'(x)=e^x,g(x)=\sqrt{x}, g'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

leave the 3 out front as a common factor, and them make a direct substitution into the formula i just wrote above

OpenStudy (anonymous):

\[3\left(e^x\sqrt{x}+e^x\times \frac{1}{2\sqrt{x}}\right)\] is the first step

OpenStudy (anonymous):

and then we multiply by 3 for each term?

OpenStudy (anonymous):

then clean up with some algebra, probably write \[3\left(\sqrt{x}e^x+\frac{e^x}{2\sqrt{x}}\right)\] you can distribute if you like, i would leave it like that

OpenStudy (anonymous):

you get more or less what i did? a direct substitution into the "product rule"

OpenStudy (anonymous):

why did you switch around the e^x's

OpenStudy (anonymous):

i get the multiplying but I'm still confused

OpenStudy (anonymous):

no real reason, just generally you see it written that way \[xe^x\] for example rather than \[e^xx\] multiplication is commutative so it makes no real difference,

OpenStudy (anonymous):

you want a more detailed explanation? i can try if you like

OpenStudy (anonymous):

well yes i understand the xe^x but why did you replace the 1 with e^x

OpenStudy (anonymous):

oooh i see your question

OpenStudy (anonymous):

at first i wrote the last term as \[e^x\times \frac{1}{2\sqrt{x}}\] then i confused you by writing it as \[\frac{e^x}{2\sqrt{x}}\]

OpenStudy (anonymous):

oh no that makes sense i get it now

OpenStudy (anonymous):

but it arithmetic with fractions is all just like \[5\times \frac{1}{3}=\frac{5}{3}\]

OpenStudy (anonymous):

you get how the product rule works? we can do another if you like

OpenStudy (anonymous):

you multiple f into g and g into f

OpenStudy (anonymous):

I have this one that I need help solving Differentiate. (Assume c is a constant.)

OpenStudy (anonymous):

f primes times g plus g prime times f

OpenStudy (anonymous):

\[z=x ^{3/2}(x+ce ^{x})\]

OpenStudy (anonymous):

would we do the exponent thing?

OpenStudy (anonymous):

for this i would multiply out first

OpenStudy (anonymous):

\[\large z=x^{\frac{5}{2}}+cxe^x\] would be my first step

OpenStudy (anonymous):

how did you get the 5/2?

OpenStudy (anonymous):

i added one to the exponent

OpenStudy (anonymous):

that is not the derivative !

OpenStudy (anonymous):

i was just multiplying out \[\large x^{\frac{3}{2}}\times x=x^{\frac{5}{2}}\]

OpenStudy (anonymous):

maybe it would be easier not to do that and go right to the product rule lets try that way

OpenStudy (anonymous):

\[z=x ^{3/2}(x+ce ^{x})\] \[z'=\frac{d}{dx}x^{3/2}(x+ce^x)+x^{3/2}\frac{d}{dx}(x+ce^x)\]

OpenStudy (anonymous):

from there we get \[z'=\frac{3}{2}x^{1/2}(x+ce^x)+x^{3/2}(1+ce^x)\] and that is all, i wouldn't change it

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw hope the steps were clear